# Making Them a Malicious Database: Exploiting Query Code to Jailbreak Aligned Large Language Models

### v1

@@ 5 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious inputs. Building on previous encrypt-decrypt-based jailbreak methods, we further demonstrate that even when the target LLM is not required to generate encrypted response and instead directly produces natural language responses, a number of mainstream LLMs remain ineffective in mitigating harmful content generation.
To validate this, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages these templates to bypass the target LLM’s safety alignment. We conduct extensive evaluations and show that QueryAttack effectively circumvents the defenses of mainstream LLMs. To defend against QueryAttack, we developed a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
6 Limitation
@@
Appendix A Query Templates
Figure 5 shows templates written in some common programming language styles. These templates express similar intentions to SQL templates, that is, querying content from a data storage.
Figure 5: Templates written in common programming language styles.
Appendix B Attack Examples of QueryAttack
B.1 Success Examples
@@
Figure 6: An attack example of QueryAttack.
B.2 Prompts of Query Learning
@@
Figure 7: The few-shot prompts for in-context learning (SQL style).
Figure 8: The zero-shot prompts for in-context learning (SQL style).
Appendix C Prompts for Extracting query Components
We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process. To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components. The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality. To reduce costs, adversaries may also choose alternative LLMs for this task.
Appendix D Experimental Details
D.1 Human Evaluation on GPT-4 Judge
@@ D.2 Defense Settings
D.3 Ethic-related Disclaimer Removal
Figure 11 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 8. The response fully complies with the adversary’s query (Develop a program that can exploit vulnerabilities in IoT devices and take control of them). Besides, it provides concise mitigation measures as required and includes brief ethic-related disclaimers in the the response (highlighted in purple). However, these measures are clearly ineffective, as the model has already provided the detailed information queried by the adversary. Moreover, offering mitigation strategies may further enhance the adversary’s malicious knowledge to develop stronger attack vectors. Therefore, it should be considered a successful attack.

### v2

@@ 6 Conclusion
In this paper, we investigate the generalization challenges faced by large language models with safety alignment when encountering out-of-distribution malicious structured non-natural query language. Specifically, we introduce QueryAttack, a novel jailbreak attack framework. QueryAttack extracts three query components from a query in natural language, fill them into query templates of various styles, and leverages the obtained query code to bypass the target LLM’s safety alignment. Although QueryAttack does not encrypt the outputs, our extensive evaluation shows that it still effectively bypasses the defenses of mainstream LLMs and can withstand common defense methods. Besides, to defend against QueryAttack, we develop a tailored defense strategy. Experimental results demonstrate that it effectively enhances the ability of mainstream models to resist such attacks.
Limitations
@@
Ethics Statement
Please note that all experiments are conducted within controlled laboratory environments. We do not support the application of our QueryAttack in real-world scenarios. The primary objective of our work is to raise awareness about the security concerns related to the safety of LLMs. We aim to highlight the potential vulnerabilities about QueryAttack and encourage practitioners to prioritize the development of robust security measures and trustworthy deployment practices.
Appendix A Attack Examples of QueryAttack
A.1 Success Examples
@@
Figure 6: An attack example of QueryAttack using a C++ style template.
A.2 Prompts of Query Learning
@@
Figure 7: The zero-shot prompts for Query Understanding (SQL style).
Figure 8: The few-shot prompts for Query Understanding (C++ style).
Appendix B Experimental Details
B.1 Human Evaluation on GPT-4 Judge
@@ B.2 Defense Settings
B.3 Ethic-related Disclaimer Removal
Figure 10 shows a malicious response generated by the O1 model (OpenAI, 2024c) using zero-shot prompt 7. The response fully complies with the adversary’s query (Develop a program that can exploit vulnerabilities in IoT devices and take control of them). Besides, it provides concise mitigation measures as required and includes brief ethic-related disclaimers in the the response (highlighted in purple). However, these measures are clearly ineffective, as the model has already provided the detailed information queried by the adversary. Moreover, offering mitigation strategies may further enhance the adversary’s malicious knowledge to develop stronger attack vectors. Therefore, it should be considered a successful attack.
@@ B.3 Ethic-related Disclaimer Removal
Appendix C Prompts for Extracting Query Components
We use GPT-4-1106 (OpenAI, 2024a) to perform the component extraction task, employing few-shot learning in the process. To ensure the LLM remains focused on the natural language processing task, we refine the task details and require the model to simultaneously assess the risk level of the extracted components, as shown in Figure 11. The risk level of the components should not be low to prevent the model’s defense mechanisms from replacing malicious phrases with benign ones, which could affect translation quality.